A ball is thrown directly up in the air from a height of 5.5 feet with an initial velocity of 65.5 ft/s. Ignoring air resistance, how long until the ball hits the ground? Use the formula h=-16t^2+65.5t+5.5, where h is the height of the ball in feet and t is the time in seconds after it is thrown. Round your answer to the nearest hundredth.
Answer by Alan3354(69443)

A ball is thrown directly up in the air from a height of 5.5 feet with an initial velocity of 65.5 ft/s. Ignoring air resistance, how long until the ball hits the ground? Use the formula h=-16t^2+65.5t+5.5, where h is the height of the ball in feet and t is the time in seconds after it is thrown. Round your answer to the nearest hundredth.
============================
h=-16t^2+65.5t+5.5 = 0
-16t^2 + 65.5t + 5.5 = 0
Solve for t.